You are given that "1+i"  is a root of the equation "z^{3}+2z^{2}+az+b=0" , where a and b are real numbers. Which of the following are true?



  1. "a=\u22126" and "b=8"
  2. "z=4"  is also a root of the equation.
  3. "z=\u22121\u2212i"  is also a root of the equation.
  4. "a=8" and "b=-6"
  5. "z=\u22124"  is also a root of the equation.
1
Expert's answer
2021-11-04T19:50:46-0400

"z^3+2z^2+az+b=0\\(1+i)^3+2(1+i)^2+a(1+i)+b=0\\1+i^3+3i(1+i)+2(1+i^2+zi)+ai+a+b=0"

"1-i+3i-3+2-2+4i+ai+a+b=0\\-2+a+b+6i+ai=0"

"-2+a+b+i(a+6)=0" (1)

therefore a +6=0

a=-6

now put this in equation 1

b=8



"z^3+2z^2+az+b=0"

now putting value of a and b

"z^3+2z^2-6z+8=0"

now lets check if -4 is a root of the equation put z=-4

"(-4)^3+2(-4)^2-6(-4)+8=0\\-64+32+24+8=0\\0=0"

  1. hence "z=\u22124"  is also a root of the equation.

(and 2,3 ,4 are incorrect)